OpenStudy (anonymous):

Factor 2x^2-19X+33=0

OpenStudy (saifoo.khan):

Sure , where are u stuck?

OpenStudy (anonymous):

I can find th fators of 33 but they do not add up or substracte to 19

OpenStudy (saifoo.khan):

that can't be factored. :p

OpenStudy (anonymous):

so what should I do about it?

OpenStudy (anonymous):

just say can't be factored?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

No problem!! ^_^

OpenStudy (saifoo.khan):

is that a dog in your photo?

OpenStudy (anonymous):

No its our biology class gunia pig

OpenStudy (anonymous):

hey I have one more question

OpenStudy (saifoo.khan):

Ohh, okay.. sure... post it.

OpenStudy (anonymous):

-x^2+8X=12

OpenStudy (anonymous):

=12+x^2+8X=0

OpenStudy (saifoo.khan):

you cannot write it that way. the x with 8 must be small.

OpenStudy (saifoo.khan):

this is also unfactorable.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so if in an equation has 2 unknown values is it still unfactorable?